Prove That:
5^31–5^29 is divisible by 100.
Format: x^y*100

Answer by Alan3354(69443)

Prove That:
5^31–5^29 is divisible by 100.
----------
5^31–5^29 = 5^29*(25 - 1)
= 24*25*5^27
= 600*5^27
= 6*5^27*100